Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2001-160_2001-166

Mission Operations Days: 2001/160 to 2001/166
June 09, 2001 00:00:00 GMT to June 15, 2001 00:00:00 GMT


The MODIS Power Supply 2 Shut Off on DOY 2001/166. (details below)

MODIS is currently in Survival Mode.


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode with SSR supersets 30 and 31 unallocated
MODIS is in Survival Mode
ComponentState(s)Comment
BlackbodyA Off; B OffPS2 Shutdown Anomaly
Calibration ElectronicsA Off; B Off (Selected)PS2 Shutdown Anomaly
Control ProcessorA Off; B OffPS2 Shutdown Anomaly
Nadir DoorUnlatched, OpenPS2 Shutdown Anomaly
Space View DoorUnlatched, OpenPS2 Shutdown Anomaly
Solar Diffuser DoorUnlatched, ClosedPS2 Shutdown Anomaly
FDDI FormatterA Off; B OffPS2 Shutdown Anomaly
FDDI PortA Off; B OffPS2 Shutdown Anomaly
FIFO Memory#1 & #2 Off; #3 & #4 OffPS2 Shutdown Anomaly
Format ProcessorA Off; B OffPS2 Shutdown Anomaly
Power Supplies#1 Off; #2 OffPS2 Shutdown Anomaly
PV VIS FPAA Off; B OffPS2 Shutdown Anomaly
PV NIR FPAA Off; B OffPS2 Shutdown Anomaly
PV SMIR FPAA Off; B OffPS2 Shutdown Anomaly
PV LWIR FPAA Off; B OffPS2 Shutdown Anomaly
PC LWIR FPAA Off; B OffPS2 Shutdown Anomaly
RC Outgas HeatersAll Off (CS, IS, OS)PS2 Shutdown Anomaly
RC LWIR FPA HeaterOffPS2 Shutdown Anomaly
RC SMIR FPA HeaterOffPS2 Shutdown Anomaly
Scan AssemblyA Off; B OffPS2 Shutdown Anomaly
SDSMA Off; B OffPS2 Shutdown Anomaly
SRCAA Off; B OffPS2 Shutdown Anomaly
PS1 Survival HeaterEnabledPS2 Shutdown Anomaly
PS2 Survival HeaterEnabledPS2 Shutdown Anomaly
Timing GeneratorA Off; B OffPS2 Shutdown Anomaly
Flight SoftwareRev BDPS2 Shutdown Anomaly
Inhibit Ids SetNonePS2 Shutdown Anomaly
TMONs Enabled68PS2 Shutdown Anomaly

This Week's Completed MODIS Activities:

Saturday, June 9, 2001

None

Sunday, June 10, 2001

Lunar Roll Maneuver (-20.0192 degree roll)
2001/161 21:35 First attitude disturbances due to configuration change for maneuver
2001/161 22:54:37 ATC Move Formatter to Night ahead by 3:23 to ensure 50/50 day/night
2001/161 22:58:54 FOT ATC commanding: Begin slew to –20.0192 degrees
2001/161 23:07:17 ATC Set Formatter to Day Rate
2001/161 23:07:19 ATC Set SCIABNORM Flag to ABNORM
2001/161 23:07:21 ATC PC DC Restore to OFF
2001/161 23:07:23 ATC PV DC Restore to OFF
2001/161 23:07:25 ATC Sector Rotation to –3072 (EA to SV and OBCs)
2001/161 23:13:55 ATC Sector Rotation to Zero (Normal)
2001/161 23:13:57 ATC PV DC Restore to ON
2001/161 23:13:59 ATC PC DC Restore to ON
2001/161 23:14:01 ATC Set SCIABNORM Flag to NORM
2001/161 23:14:03 ATC Set Formatter to Night Rate
2001/161 23:14:35 FOT ATC commanding: Return to nominal attitude
2001/161 23:24 Final attitude disturbances due to configuration change for maneuver
2001/161 23:50:52 ATC Move Formatter to Day later by 3.23 to ensure 50/50 day/night

Monday, June 11, 2001

None

Tuesday, June 12, 2001

None

Wednesday, June 13, 2001

2001/164 15:53:07-16:04:35 ATC OA15: SD/SDSM Open
2001/164 16:44:21-17:17:54 ATC OA19: SRCA Full Radiometric
2001/164 17:32:00-17:43:28 ATC OA16: SD/SDSM Screened
2001/164 19:27:39-19:30:10 ATC OA27: PV ECAL
2001/164 21:05:31-21:10:05 ATC OA28: PC ECAL

Thursday, June 14, 2001

None


Friday, June 15, 2001

2001/166 05:16:43-05:23:34 Real-Time MOD_ALL_OFF Sent in response to anomalous telemetry.


This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:


No events will be scheduled until PS2 Shut Off Anomaly is resolved.


Saturday, June 15, 2001

None

Sunday, June 16, 2001

None

Monday, June 17, 2001

None

Tuesday, June 18, 2001

None

Wednesday, June 19, 2001

None

Thursday, June 20, 2001

None

Friday, June 21, 2001

None

Maneuvers:

The next Drag Make Up maneuver will occur on DOY 178. It will be a 16 second burn beginning at approximately 19:05.

The next MODIS Lunar Roll Calibration is scheduled for DOY 191 from 15:07:42 to 15:14:12. This roll has the obvious prerequisite that MODIS must successfully from its Power Supply 2 Anomaly (described below).

MODIS Anomalies:

On June 15th, 2001, at 03:56:08 GMT, MODIS experienced an anomaly.

The Flight Operations Team paged the on-call MODIS engineer reporting a trip of TMON #68, the MODIS double FIFO write TMON. In addition, many red limit violations were reported. Inspection of telemetry showed peculiar states including a CP mode of INVALID and the MODIS model as ENGINEERING. With all telemetry suspect, the MODIS IOT commanded the instrument to low-power mode (using MOD_ALL_OFF) at 05:16:43.

Investigation is still pending, but below is the current understanding of the event:

Continued telemetry analysis showed that the TMON itself did not trip, but rather, it went into an error state because it was receiving stale telemetry. At the anomaly time, all telemetry “pegged out”, and the telemetry headers stopped incrementing (thus explaining the error reported by the TMON).

Science data stopped being recorded on the SSR at the same time the housekeeping telemetry “pegged out”. This suggests that it was NOT a 1553 Bus communication problem between MODIS and the spacecraft, but rather something that would affect both housekeeping and science data.

At the time of the anomaly, the EPS feed current to MODIS (an external spacecraft point) dropped from ~1.5 Amps to 0.07 Amps suggesting that all MODIS components, with the exception of the power supply, powered down at that time.

An Anomaly review/recovery team has been assembled and has begun holding daily meetings to determine the cause, and recovery path from this anomaly.

General Instrument Comments:

MODIS is currently powered off in Survival Mode. The Survival Heaters are cycling as expected. Return to an operational mode is currently pending the outcome of the MODIS Power Supply 2 Anomaly.

MODIS Telemetry Trends:

Telemetry is trending nominally. The amount of power required by the focal plane heater to maintain 83.0K is still slowly decreasing, although the spacecraft is currently near the minimum absolute solar beta angle for the year.

Voltage, current and temperature trending over the previous month, in addition to selected past time periods is currently underway to support the resolution of the MODIS Power Supply 2 Shut Down Anomaly.

Non-MODIS Significant Events:

MOPITT Anomaly:

The MOPITT team is still working to recover their instrument to science operations.

Limited Life Item Status:

SRCA 10W Lamp #1: 186.1 of 500 hours
SRCA 10W Lamp #2: 143.0 of 500 hours
SRCA 10W Lamp #3: 152.2 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 557.6 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1633 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements